# SoundMind: RL-Incentivized Logic Reasoning for Audio-Language Models

### v1

@@ 5.1 Setup
We fine-tune the Qwen2.5-Omni-7B model on the ALR dataset using the proposed SoundMind algorithm. All experiments were conducted under a consistent hardware environment, consisting of an Intel(R) Xeon(R) Platinum 8468 CPU and 8 NVIDIA H800 GPUs, each equipped with 80 GB of memory. The weighting coefficients for the reward components were set as follows: $\lambda_{1}=1.0$, $\lambda_{2}=0.5$, $\lambda_{3}=2.0$, $\lambda_{4}=1.0$, and $\lambda_{5}=0.75$. The training procedure was executed for 50,000 steps. All other hyperparameter settings followed those used in Logic-RL Xie et al. (2025a).
To consider the dual-modality nature of both input and output in the ALR dataset, we conduct experiments across three settings: 1 Table 4: audio-only input with text-only reasoning output. We select five multimodal LLMs as baselines: MiniCPM-o (Yao et al., 2024), Gemini-Pro-V1.5 (Team et al., 2024), Baichuan-Omni-1.5 (Li et al., 2025), Qwen2-Audio (Chu et al., 2024), and Qwen2.5-Omni-7B (Xu et al., 2025). 2 Table 5: text-only input with audio-only reasoning output. Since there are few bimodal LLMs with a similar number of parameters to Qwen2.5-Omni-7B, we use Qwen2.5-Omni-7B as the baseline. 3 Table 6: audio-only input with audio-only reasoning output.
@@ 5.2 Dataset Analysis
Table 3 presents statistics of the ALR dataset, highlighting its scale, balance, and linguistic richness across training, validation, and test splits. The dataset comprises a total of 6,446 samples, with a relatively balanced distribution between entailed and not-entailed instances, ensuring fair coverage of both classes for entailment prediction.
The average length of user content is approximately 160–180 tokens across all splits, while the generated CoT responses are significantly longer, averaging over 1,400 tokens. This discrepancy underscores the dataset’s emphasis on step-by-step, explainable reasoning, which is crucial for supporting interpretability in audio-based reasoning tasks.
In terms of audio representation, the user content audio averages around one minute in duration, whereas the CoT reasoning speech spans nearly 10 minutes per sample. This extended audio format provides an excellent benchmark for evaluating models’ long-form audio understanding and reasoning capabilities.
As shown in Figure 4, the ALR dataset exhibits a wide range of duration distributions that reflect critical considerations for audio-logical reasoning research. Audio segments predominantly span 3–12 minutes (with 28.6% of training samples in the 540–720s range), balancing sufficient contextual information with manageable computational complexity. The design deliberately avoids shorter clips lacking substantive reasoning content.
The dataset demonstrates a modest class imbalance (44.9% entailed vs. 55.1% not-entailed overall), mirroring real-world reasoning scenarios. This imbalance is particularly pronounced in mid-range durations (240–420s), where audio complexity appears to challenge entailment determination. Notably, the 480–540s bin in the training set shows an inverse distribution (56.9% entailed), possibly reflecting unique acoustic features or annotation artifacts.
Strategic split allocation maintains consistent entailment ratios across training (42.8%), test (54.9%), and validation (56.3%) sets, ensuring reliable evaluation while preserving inherent complexity. Proportional representation across all duration bins in each split helps prevent duration-based bias. The near-balanced nature of the test set (360 entailed vs. 296 not-entailed) supports rigorous and fair model assessment.
Overall, this distribution profile directly supports the dataset’s objective of advancing bimodal reasoning by providing diverse yet controlled audio-text interaction scenarios. These settings challenge models to develop genuine cross-modal understanding, rather than relying on duration-specific pattern recognition.
@@ 5.3 Results and Analysis
Model Accuracy (%$\uparrow$
@@
Qwen2.5-Omni-7B-RL (Our) 81.40
Table 4: Test accuracy (%) of different models on the ALR benchmark for audio-to-text logical reasoning (audio input, text output).
SoundMind achieves state-of-the-art performance on the audio-to-text reasoning task. Table 4 presents results for generating textual reasoning outputs from audio-only inputs. Our Qwen2.5-Omni-7B-RL model achieves an accuracy of 81.40%, outperforming all evaluated baselines. Compared to its supervised counterpart Qwen2.5-Omni-7B (77.59%), the reinforcement-tuned variant achieves a notable absolute gain of 3.81%. This indicates that the SoundMind reward framework significantly improves the model’s ability to generate accurate, text-based logical conclusions based on audio inputs. Among non-RL baselines, Gemini-Pro-V1.5 achieves the second-best performance with 74.54%, followed by MiniCPM-o (73.17%), and Baichuan-Omni-1.5 (70.58%). Qwen2-Audio lags behind with 58.23%, highlighting that generic multimodal alignment is insufficient for reasoning-heavy tasks without dedicated optimization. Qwen2.5-Omni-7B-RL improves upon Gemini-Pro by a margin of 6.86%, and surpasses MiniCPM-o by 8.23%, establishing a clear lead across models of comparable or larger capacity. These results validate that our reinforcement learning setup improves reasoning accuracy and generalizes effectively across audio entailment tasks, driven by structured reward signals for answer format, answer correctness, and reasoning length.
SoundMind enhances logical structuring in text-to-audio reasoning despite modality shift. In Table 5, we evaluate the model’s ability to generate speech-based reasoning from text-only prompts. The RL-enhanced model achieves 83.84% accuracy, representing a 3.05% improvement over the supervised baseline. This result highlights that even when the input lacks audio grounding, the reward structure guides the model toward more structured and consistent reasoning in speech. The increased reasoning depth, induced by chain-of-thought reinforcement signals, indicates that SoundMind is capable of refining response quality across modalities—not only in interpreting audio, but also in generating semantically coherent outputs in speech form.
Model WER (%$\downarrow$ Acc. (%$\uparrow$
@@
Qwen2.5-Omni-7B-RL (Our) 6.99 83.84
@@
In fully speech-based settings, SoundMind generalizes reasoning without textual anchors. Table 6 investigates the most demanding configuration: generating spoken reasoning from audio-only inputs, with no access to textual content during inference. In this dual-audio scenario, our model maintains a strong performance gain, raising accuracy from 77.59% to 81.40%. This demonstrates that SoundMind can internalize abstract logical dependencies in audio alone—without relying on textual shortcuts. Such robustness is especially valuable for real-world applications involving spoken dialogue systems, voice-based tutoring agents, or accessibility tools that require full audio-audio reasoning capabilities.
Across all tasks, SoundMind consistently improves logical accuracy while maintaining acceptable speech quality. Our Qwen2.5-Omni-7B-RL model outperforms all baselines in reasoning accuracy across the three evaluated input-output modality pairs. These results confirm that reward-guided reinforcement learning can strengthen logical inference capabilities in audio-language models. While we observe a moderate rise in Word Error Rate (WER) in speech-based outputs, this trade-off is expected given the model’s increased focus on semantic correctness and structured explanation. The observed WER remains within acceptable bounds and does not hinder interpretability. Future work can explore fluency-aware RL objectives or dual-path reward balancing to further harmonize reasoning fidelity and audio naturalness.
Model WER (%$\downarrow$ Acc. (%$\uparrow$
@@
Qwen2.5-Omni-7B-RL (Our) 8.95 81.40
[…]
@@ 6 Conclusion
We introduce SoundMind, a novel rule-based reinforcement learning framework that empowers large-scale audio-language models with advanced logical reasoning capabilities across both audio and textual modalities. To enable such training, we build the Audio Logical Reasoning (ALR) dataset, a dual-modality benchmark comprising 6,446 high-quality samples annotated with chain-of-thought reasoning in both audio and text forms. Experimental results demonstrate that our method significantly improves performance and establishes new state-of-the-art results on the ALR benchmark, consistently outperforming strong baselines across three reasoning settings: text-to-audio, audio-to-text, and audio-to-audio. We hope this work can serve as a foundation for future research in logic-oriented large-scale audio-language modeling and in reinforcement-driven multimodal reasoning.
@@
While our approach demonstrates strong performance, certain aspects warrant further exploration. In particular, the Word Error Rate (WER) in audio generation remains competitive, but still leaves room for improving reasoning fidelity during speech synthesis. Addressing this limitation could lead to more robust, adaptive, and human-aligned audio reasoning systems in future work.

### v2

@@ 5.1 Setup
We fine-tune the Qwen2.5-Omni-7B model on the SoundMind dataset using the proposed SoundMind-RL algorithm. All experiments were conducted under a consistent hardware environment, consisting of an Intel(R) Xeon(R) Platinum 8468 CPU and 8 NVIDIA H800 GPUs, each equipped with 80 GB of memory. The weighting coefficients for the reward components were set as follows: $\lambda_{1}=1.0$, $\lambda_{2}=0.5$, $\lambda_{3}=2.0$, $\lambda_{4}=1.0$, and $\lambda_{5}=0.75$. The training procedure was executed for 50,000 steps. All other hyperparameter settings followed those used in Logic-RL Xie et al. (2025a).
To fully exploit the availability of both text and audio modalities in the SoundMind dataset, we evaluate our approach under three input–output configurations: 1 Table 4: audio-only input with text-based reasoning output, benchmarked against five strong multimodal LLMs: MiniCPM-o (Yao et al., 2024), Gemini-Pro-V1.5 (Team et al., 2024), Baichuan-Omni-1.5 (Li et al., 2025), Qwen2-Audio (Chu et al., 2024), and Qwen2.5-Omni-7B (Xu et al., 2025). 2 Table 5: text-only input with audio-based reasoning output. Due to the lack of comparable open-source multimodal models, we use Qwen2.5-Omni-7B as the reference baseline. 3 Table 6: audio-only input with audio-based reasoning output. Together, these three configurations allow us to comprehensively assess SoundMind-RL across text and audio modalities, covering fully cross-modal reasoning scenarios.
@@ 5.2 Dataset Analysis
Table 3 summarizes key statistics of the SoundMind dataset, clearly illustrating its overall scale, class balance, and linguistic diversity across training, validation, and test splits. The corpus contains 6,446 samples with a near-balanced distribution of entailed and not-entailed instances, thereby ensuring representative coverage of both classes for robust entailment prediction.
The user content averages 160–180 tokens across all splits, whereas the generated CoT responses are much longer, exceeding 1,400 tokens on average. This substantial length gap reflects the dataset’s focus on detailed, step-by-step reasoning, providing rich supervision that encourages faithful and transparent reasoning in audio-based tasks.
In terms of audio representation, the user content speech averages roughly one minute per sample, whereas the CoT reasoning audio extends to nearly ten minutes on average. This substantial duration provides a challenging and realistic benchmark for assessing models’ ability to perform long-form audio comprehension and sustained logical reasoning.
As shown in Figure 4, the SoundMind dataset covers a broad range of audio durations, capturing important variability for developing and evaluating models for audio-based logical reasoning. Most segments fall within 3–12 minutes, with 28.6% of training samples concentrated in the 540–720s range, which provides ample contextual information while keeping computational cost tractable. Notably, the dataset deliberately excludes very short clips to ensure that each sample contains sufficiently meaningful reasoning content.
The dataset exhibits a class distribution of 44.9% entailed and 55.1% not-entailed instances overall, closely resembling patterns in many real-world reasoning scenarios. This skew becomes more pronounced in the mid-range durations (240–420s), suggesting that longer and acoustically richer segments may introduce additional difficulty for reliable entailment classification. Notably, the 480–540s range in the training split shows a reversed distribution, with 56.9% entailed samples, potentially reflecting distinctive acoustic patterns or systematic properties of the original source data.
The SoundMind dataset is split to maintain consistent entailment ratios across training (42.8%), test (54.9%), and validation (56.3%) sets, thereby enabling reliable evaluation while preserving the inherent task complexity. Each split carefully preserves proportional coverage across all duration bins, effectively minimizing potential duration-related bias and ensuring representative diversity. The test set is nearly balanced (360 entailed vs. 296 not-entailed), providing a solid basis for rigorous and fair model assessment.
Overall, this distribution profile aligns with the dataset’s goal of advancing audio–language reasoning research by offering diverse but systematically controlled interaction scenarios. Such a design compels models to acquire robust cross-modal reasoning skills, discouraging reliance on superficial duration-based correlations and encouraging deeper semantic understanding.
@@ 5.3 Results and Analysis
5.3.1 Audio-to-Text Reasoning
[…]
@@ 5.3.1 Audio-to-Text Reasoning
Qwen2.5-Omni-7B (SoundMind-RL) 81.40
Table 4: Accuracy (%) of evaluated models on the SoundMind benchmark under the audio-to-text configuration. In this setting, models receive audio-only inputs and produce textual outputs containing the final answer. The table compares several strong multimodal LLM baselines with the reinforcement-tuned Qwen2.5-Omni-7B (SoundMind-RL).
Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves state-of-the-art performance on the challenging audio-to-text reasoning task. Table 4 reports results for generating textual outputs from audio-only inputs. Our model attains an accuracy of 81.40%, clearly outperforming all evaluated baselines. Compared to its counterpart Qwen2.5-Omni-7B (77.59%), the reinforcement-tuned variant yields an absolute improvement of 3.81%, clearly demonstrating that the SoundMind reward framework substantially enhances the model’s ability to produce accurate text-based logical conclusions from speech audio input.
Among the remaining baseline models, Gemini-Pro-V1.5 reaches 74.54%, followed closely by MiniCPM-o (73.17%) and Baichuan-Omni-1.5 (70.58%). Qwen2-Audio performs the weakest, achieving only 58.23%, clearly underscoring that generic multimodal alignment is insufficient for reasoning-intensive tasks without specialized optimization. Overall, Qwen2.5-Omni-7B (SoundMind-RL) further improves upon Gemini-Pro by 6.86% and surpasses MiniCPM-o by 8.23%, establishing a consistently clear lead across models of comparable or larger capacity. These findings strongly confirm that our reinforcement learning approach strengthens reasoning accuracy and generalizes effectively across audio reasoning tasks.
5.3.2 Text-to-Audio Reasoning
[…]
@@ 5.3.2 Text-to-Audio Reasoning
Qwen2.5-Omni-7B (SoundMind-RL) 6.99 83.84
Table 5: Performance of Qwen2.5-Omni-7B and its SoundMind-RL fine-tuned variant on the SoundMind benchmark under the text-to-audio setting. In this configuration, models receive text-only inputs and generate audio outputs. Results are reported using Word Error Rate (WER) to assess speech fidelity and accuracy (%) to measure reasoning correctness.
SoundMind-RL significantly improves reasoning performance in the text-to-audio setting. As shown in Table 5, Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves an accuracy of 83.84%, yielding a 3.05% improvement over the Qwen2.5-Omni-7B baseline. These results clearly indicate that the reward mechanism effectively encourages the model to produce coherent and consistent spoken reasoning, even without audio input. The observed accuracy gain is accompanied by a higher WER (6.99% vs. 2.18%), which can be attributed to the generation of longer and more detailed reasoning sequences under reinforcement learning, thereby increasing the chance of recognition errors while preserving semantic correctness.
5.3.3 Audio-to-Audio Reasoning
@@
Qwen2.5-Omni-7B (SoundMind-RL) 8.95 81.40
Table 6: Performance of Qwen2.5-Omni-7B and its SoundMind-RL fine-tuned variant on the SoundMind benchmark under the audio-to-audio reasoning setting. In this configuration, models receive audio-only inputs and generate audio reasoning outputs.
In the fully audio-based setting, our SoundMind-RL framework delivers notable performance gains under the most challenging condition. Table 6 presents results where models must derive and verbalize reasoning solely from acoustic input. Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves 81.40% accuracy, improving substantially over the baseline (77.59%) and showing that the model can extract task-relevant logical dependencies directly from audio signals. The result is accompanied by a higher WER (8.95% vs. 2.23%), which we attribute to the generation of longer, more elaborate reasoning sequences and the added difficulty of learning from raw acoustic input. Our analysis attributes this to the absence of explicit fluency constraints during reinforcement learning, which instead prioritizes structured reasoning and semantic correctness.
Takeaway. Across all evaluated settings, SoundMind-RL consistently enhances logical reasoning performance for audio–language models. Qwen2.5-Omni-7B fine-tuned with SoundMind-RL achieves substantial gains across audio-to-text, text-to-audio, and audio-to-audio configurations, outperforming all baseline models. These results demonstrate that reward-guided reinforcement learning can reliably improve multimodal reasoning accuracy by encouraging faithful answer formatting, factual correctness, and sufficiently detailed reasoning traces. For speech-based outputs, we observe a moderate increase in WER relative to the baseline, which correlates with the production of longer, information-rich reasoning segments and suggests that the model prioritizes semantic completeness over surface fluency. Future work can explore fluency-aware or prosody-sensitive reward components to better balance reasoning fidelity with the naturalness of generated speech.
@@ 6 Conclusion
We introduce SoundMind-RL, a novel rule-based reinforcement learning framework that empowers large-scale audio-language models with advanced logical reasoning capabilities across both audio and textual modalities. To support such training, we construct SoundMind, a dataset for the task of audio logical reasoning, containing 6,446 carefully curated audio–text pairs, each annotated with modality-specific chain-of-thought reasoning. Experimental results demonstrate that our method significantly improves performance and establishes state-of-the-art results on the SoundMind benchmark, consistently outperforming strong baselines across three reasoning settings: text-to-audio, audio-to-text, and audio-to-audio. We hope this work provides a useful resource and framework for advancing research in reasoning-oriented audio–language modeling and reinforcement-based multimodal learning.
@@
While this study shows promising results, several limitations remain. (1) The rule-based reward design enforces reasoning format consistency but may introduce rigidity, and its generalization to more diverse or open-ended tasks remains to be examined. (2) The dataset relies on synthetic speech and automatically generated chain-of-thought annotations; despite careful curation, subtle artifacts or biases may still be present and could influence model behavior. (3) The increase in Word Error Rate (WER) observed in audio generation suggests a possible trade-off between reasoning depth and fluency, motivating further work on prosody modeling and decoding strategies. We leave these challenges for future work, aiming to advance the development of audio reasoning systems that are more robust, generalizable, and aligned with human needs.
@@ Acknowledgment
Appendix A Baselines
  • MiniCPM-o Yao et al. (2024): An open-source multimodal language model supporting vision, speech, and real-time streaming. It emphasizes efficiency and deployability, making it suitable for interactive multimodal understanding and generation across a variety of general-purpose and resource-constrained scenarios.
  • Gemini-Pro-V1.5 Team et al. (2024): A multimodal model developed by Google that processes text, image, audio, and video inputs in a unified way. It is designed to handle extended context reasoning and cross-modal alignment, enabling coherent understanding and response in complex, multi-step tasks across modalities.
  • Baichuan-Omni-1.5 Li et al. (2025): A multimodal model that combines text, vision, and audio understanding with audio generation. It aims to enable coherent cross-modal interaction and supports a wide range of multimodal and multitask applications in interactive settings.
  • Qwen2-Audio Chu et al. (2024): A speech language model that accepts both audio and text as input for natural voice-based interaction. It can perform detailed audio analysis, transcription, and multilingual understanding, serving as a general-purpose audio–text reasoning system.
  • Qwen2.5-Omni-7B Xu et al. (2025): A multimodal model that handles text, image, audio, and video inputs and generates both text and speech outputs. It supports streaming interactions and aims to provide robust multimodal reasoning and natural communication in real-time applications.
Appendix B Text-to-Speech Model
Constructing a fully human-recorded version of the audio logical reasoning dataset—covering thousands of user prompts, chain-of-thought reasoning traces, and final answers in both input and output modalities—would require extensive scripting, careful speaker coordination, and thousands of hours of professional narration. Such a large-scale recording effort is prohibitively expensive and logistically infeasible given the diversity and duration of reasoning sequences we target. Consistent with established practice in speech and multimodal learning research Jia et al. (2022); Zevallos (2022); Regmi et al. (2019); Cheng et al. (2024); Bartelds et al. (2023); Zhong et al. (2022), we synthesize the audio modality using a high-fidelity text-to-speech (TTS) system. This strategy enables scalable and reproducible generation of paired input–output audio segments, while preserving fine control over speaking rate, prosody, and acoustic quality—key factors for training models to perform multi-minute reasoning without performance degradation.
To ensure that the synthesized audio faithfully reflects the complexity of logical reasoning, we conducted a systematic evaluation of several state-of-the-art open-source TTS systems. We compared MegaTTS-3 Jiang et al. (2025), WhisperSpeech OpenAI (2024), Spark-TTS Wang et al. (2025), and Zonos Zyphra (2024) across multiple dimensions, including perceived naturalness, prosody stability over long contexts, noise robustness, and intelligibility of extended reasoning chains. Each candidate system was benchmarked by synthesizing representative samples from our dataset and evaluated via multi-rater human listening studies. MegaTTS-3 consistently achieved the highest mean opinion scores, particularly in naturalness and prosodic consistency, qualities that we consider essential for faithfully preserving the chain-of-thought signal.
Configuration $\lambda_{1}$ (TextFmt) $\lambda_{2}$ (AudioFmt) $\lambda_{3}$ (Answer) $\lambda_{4}$ (TextLen) $\lambda_{5}$ (AudioLen) Accuracy (%) $\Delta$ vs. full (pp)
(1) w/o audio rewards 1.0 0 2.0 1.0 0 70.82 -10.58
(2) w/o text rewards 0 1.0 2.0 0 1.0 48.84 -32.56
(3) w/o answer reward 1.0 0.5 0 1.0 0.75 60.24 -21.16
(4) full (Ours) 1.0 0.5 2.0 1.0 0.75 81.40
Table 7: Ablation of the composite reward in SoundMind-RL, evaluated on the test set. $\lambda_{1}/\lambda_{2}$ enforce text/audio format compliance, $\lambda_{3}$ supervises answer correctness, and $\lambda_{4}/\lambda_{5}$ encourage reasoning length in text/audio space. Accuracy is reported in %, and $\Delta$ indicates percentage-point ($p$) difference relative to the full model.
On the basis of these findings, we select MegaTTS-3 as the synthesis engine for both user prompts and model reasoning traces. Its ability to produce long, coherent, and acoustically stable speech makes it well suited for the extended reasoning sequences characteristic of the ALR task. The resulting corpus provides a controlled yet realistic training signal that aligns with our objective of enabling robust, audio-native logical reasoning in large audio-language models. We further acknowledge the open-source speech community for providing high-quality TTS systems, which have been instrumental in enabling reproducible, large-scale construction of the audio datasets.
Appendix C Potential Applications
Reasoning-capable ALMs can enable applications that demand explainability, transparency, and natural interaction. We highlight key domains where audio–native reasoning offers clear value and outline open challenges for future research.
Education and Tutoring.
Spoken step-by-step explanations can help learners understand complex problems by exposing intermediate reasoning rather than presenting only a final answer Wang et al. (2020). This is particularly relevant for subjects such as logic, mathematics, and music, where reasoning over multimodal signals is essential for mastery. Recent studies on music-related question answering and multimodal reasoning further highlight the need for models that can integrate audio and symbolic information to support learning and feedback You et al. (2025); Diao et al. (2025).
Accessibility and Screen-Free Interaction.
For users with visual impairments, hands-busy scenarios, or clinical settings, verbal reasoning offers an inclusive interface to access complex information. Replayable and summarizable reasoning can support comprehension and patient-centered decisions, consistent with efforts on trustworthy medical response generation Li et al. (2024) and healthcare voice assistants Zhan et al. (2024). Ensuring natural prosody and coherence over multi-minute speech remains crucial for high-stakes applications.
Conversational Agents and Voice Assistants.
In applications such as troubleshooting, procedural guidance, and eligibility checking, users often ask “why” in addition to “what.” Models capable of structured, spoken justifications could reduce repeated follow-ups and improve user trust.
Meetings, Lectures, and Knowledge Capture.
Beyond transcription, audio–native reasoning models could verbalize intermediate inferences, summarize discussions, and surface entailment relations in real time, turning speech streams into structured knowledge. Robustness to overlapping speech, noise, and speaker variability remains a key challenge, as exemplified by the AMI Meeting Corpus Carletta et al. (2005).
Appendix D Ablation Analysis
Quantitative impact.
Table 7 reports the results of ablating individual reward components. Removing text-format/length supervision (Row 2) causes the most severe degradation, with accuracy dropping to 48.84% (-32.56% relative to the full model). This sharp decline underscores that structural guidance on the text side is essential for producing well-formed and logically verifiable reasoning traces. Excluding the answer-correctness reward (Row 3) reduces accuracy to 60.24% (-21.16%), indicating that explicit factual supervision is necessary to ensure generated reasoning converges to the correct entailment decision rather than drifting into irrelevant but superficially coherent explanations. Removing audio-format/length rewards (Row 1) still lowers accuracy to 70.82% (-10.58%), suggesting that modality-consistent formatting and duration constraints regularize the shared policy and enhance robustness. Overall, these results demonstrate that each reward term contributes meaningfully to model performance, and that their combination is crucial for maximizing reasoning accuracy.
Key takeaways.
Table 7 highlights three main takeaways: (1) Text-side format and length supervision is the dominant factor for producing reliable, verifiable reasoning traces. (2) Answer-correctness reward is essential to align reasoning with ground truth labels and convert fluent explanations into correct decisions. (3) Audio-side rewards serve as powerful regularizers, improving policy robustness and benefiting reasoning quality across modalities.